OpenStudy (anonymous):

Multiply or divide each expression below and simplify the result. Assume the denominator is not equal to zero...

OpenStudy (anonymous):

\[\frac{ 10x^2-11x+3 }{ x^2-6x-40 }*\frac{ x^2+11x+28 }{ 2x^-x }\div \frac{ x+7 }{ 2x^2-20x }\]

OpenStudy (anonymous):

I got ...\[\frac{ 2(-2x+1 )}{ 2x-1 }\] but a math calculator got ...\[10x-6\] I want to know where I went wrong.

OpenStudy (jebonna):

@paki

OpenStudy (anonymous):

@paki

OpenStudy (jebonna):